OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –5) and is perpendicular to the line y = 3x – 6?

OpenStudy (experimentx):

y+5 = -1/3*(x-3)

OpenStudy (anonymous):

any line perp to a given line then m1*m2=-1 m1=slope of known line m2=slope of unknown line then find slope and put it in y-y1=m(x-x1)

OpenStudy (anonymous):

Here are the choices y = –3x – 4 y = –3x + 4 y = x – 4 y = x + 4

OpenStudy (anonymous):

try it yourself i gave you the hint!!!!